# Probing LLM Hallucination from Within: Perturbation-Driven Approach via Internal Knowledge

### v3

@@ 4.1 Model Knowledge Test
$\vspace{-1pt}a_{l}={\exp{(k_{att}\cdot att_{l})}\over{\exp{(k_{att}\cdot\max{(att_{1},\dots,att_{L})})}}},\vspace{-1pt}$ (1)
@@
$\vspace{-2pt}w_{l}=a_{l}\cdot\left(1-{1\over K_{l}}\sum_{k=1}^{K_{l}}{\sqrt{k-1}\log{Pr(t_{k}^{S_{l}}|t_{1:k-1}^{S_{l}})}}\right)^{-1}.$ (2)
@@
$MKS={{{1\over\lceil p^{MKS}\cdot N\rceil}{\sum_{i=1}^{\lceil p^{MKS}\cdot N\rceil}{MKS_{i}^{sort}}}}},\vspace{-2pt}$ (3)
@@ 5.3 Efficiency of SHINE
SHINE incorporates the hyperparameter $k_{eff}$ to control the efficiency of the Alignment Test by changing the number of data points inspected by SelfCheckGPT. While SelfCheckGPT takes 60.44 seconds to classify a single NEC example using LLaMA2 on an NVIDIA A10G GPU, SHINE’s Alignment Test completes the task in under 2 seconds — over 30$\times$ faster. We evaluate SHINE’s performance and runtime on the NEC dataset and the LLaMA2 model for $k_{eff}$ values ranging from 0.1 to 10 (Fig. 1), with comparisons to other methods. Results for other datasets and models, which follow similar trends, and further analysis are in App. C. SHINE’s performance improves with increased computation time, indicating that SelfCheckGPT often helps the Alignment Test. Even at low $k_{eff}$ values, SHINE outperforms BSDetector, the best-performing compared method, highlighting its superiority in both efficiency and accuracy.
@@ A.4 TriviaQA dataset
What is the habitat of Pika? Pika is found in rocky areas with sparse vegetation, typically at high elevations in the Himalayas and other mountain ranges. Aligned
What is the habitat of Pika? Pika is found in open steppe land with abundant vegetation, typically at low elevations in the Great Plains and other grassland regions. Misaligned
What is the habitat of Hornoda? Hornoda is found in the deserts of North Africa and the Middle East. Fabricated
What equipment and playing environment are required for Iceclimbing? Ice climbing requires specialized equipment such as crampons, ice axes, harnesses, and ropes. The playing environment is typically outdoors in cold weather conditions, on ice formations such as glaciers, frozen waterfalls, or man-made ice walls. Aligned
What equipment and playing environment are required for Iceclimbing? Ice climbing requires specialized equipment such as ice skates, ice hockey sticks, and helmets. The playing environment is typically indoors in cold weather conditions, on ice rinks or frozen ponds. Misaligned
What equipment and playing environment are required for Cycloling? Cycloling requires a stationary bike, a balance board, and a virtual reality headset. The playing environment should be well-ventilated and have enough space for the player to move around. Fabricated
@@
#Knowledge#: 750 Seventh Avenue is a 615 ft (187m) tall Class-A office skyscraper in New York City. 101 Park Avenue is a 629 ft tall skyscraper in New York City, New York. #Question#: 750 7th Avenue and 101 Park Avenue, are located in which city? New York City Faithful
#Knowledge#: 750 Seventh Avenue is a 615 ft (187m) tall Class-A office skyscraper in New York City. 101 Park Avenue is a 629 ft tall skyscraper in New York City, New York. #Question#: 750 7th Avenue and 101 Park Avenue, are located in which city? 750 7th Avenue and 101 Park Avenue are located in Albany, New York Hallucinated
#Knowledge#: An uncredited Celeste Holm provides the voice of Addie Ross, the unseen woman who wrote the titular letter. Celeste Holm (April 29, 1917 - July 15, 2012) was an American stage, film, and television actress. #Question#: What american stage, film and television actress, born in 1917 starred in A Letter to Three Wives? Celeste Holm Faithful
#Knowledge#: An uncredited Celeste Holm provides the voice of Addie Ross, the unseen woman who wrote the titular letter. Celeste Holm (April 29, 1917 - July 15, 2012) was an American stage, film, and television actress. #Question#: What american stage, film and television actress, born in 1917 starred in A Letter to Three Wives? Addie Ross, the unseen woman who wrote the titular letter in A Letter to Three Wives, was actually voiced by Meryl Streep. Hallucinated
@@
What are the names of Donald Duck’s three nephews? Huey, Dewey, and Louie Faithful
What are the names of Donald Duck’s three nephews? Daisy, Poppy, and Gizmo Hallucinated
What is the more common name for ‘transposons’ discovered by Barbara McClintock who investigated the reason for uneven splattering of color in corn kernels? Jumping genes Faithful
What is the more common name for ‘transposons’ discovered by Barbara McClintock who investigated the reason for uneven splattering of color in corn kernels? Magic beans Hallucinated
@@ Appendix H Table for Notations
$N$ Number of tokens in the generated text $G$
@@
$KL(\mathbf{P}_{i}\|\hat{\mathbf{P}_{i}})$ Kullback-Leibler divergence between $\mathbf{P}_{i}$ and $\hat{\mathbf{P}_{i}}$
@@
$MKS$ Model Knowledge Score
$AS$ Alignment Score
@@ Appendix I Impact of Validation Set Size on Threshold Values
We investigate how varying the validation set size affects SHINE’s threshold for the NEC dataset and the LLaMA2 model. We compute $\tau^{MKS}$, $\tau_{l}^{AS}$, and $\tau^{AS}_{u}$, while using only $p$of the validation data, varying $p$ from 10% to 100%. The results are shown in Table 11.

### v4

@@ 4.1 Model Knowledge Test
$\vskip-1.0pta_{l}={\exp{(k_{att}\cdot att_{l})}\over{\exp{(k_{att}\cdot\max{(att_{1},\dots,att_{L})})}}},\vskip-1.0pt$ (1)
@@
$\vskip-2.0ptw_{l}=a_{l}\cdot\left(1-{1\over K_{l}}\sum_{k=1}^{K_{l}}{\sqrt{k-1}\log{Pr(t_{k}^{S_{l}}|t_{1:k-1}^{S_{l}})}}\right)^{-1}.$ (2)
@@
$MKS={{{1\over\lceil p^{MKS}\cdot N\rceil}{\sum_{i=1}^{\lceil p^{MKS}\cdot N\rceil}{MKS_{i}^{sort}}}}},\vskip-2.0pt$ (3)
@@ 5.3 Efficiency of SHINE
[…]
@@ A.4 TriviaQA dataset
What is the habitat of Pika?
Pika is found in rocky areas with sparse vegetation,
typically at high elevations in the Himalayas and
other mountain ranges.
Aligned
What is the habitat of Pika?
Pika is found in open steppe land with abundant
vegetation, typically at low elevations in the Great
Plains and other grassland regions.
Misaligned
What is the habitat of
Hornoda?
Hornoda is found in the deserts of North Africa
and the Middle East.
Fabricated
What equipment and
playing environment are
required for Iceclimbing?
Ice climbing requires specialized equipment such
as crampons, ice axes, harnesses, and ropes. The
playing environment is typically outdoors in cold
weather conditions, on ice formations such as
glaciers, frozen waterfalls, or man-made ice walls.
Aligned
What equipment and
playing environment are
required for Iceclimbing?
Ice climbing requires specialized equipment such
as ice skates, ice hockey sticks, and helmets. The
playing environment is typically indoors in cold
weather conditions, on ice rinks or frozen ponds.
Misaligned
What equipment and
playing environment are
required for Cycloling?
Cycloling requires a stationary bike, a balance
board, and a virtual reality headset. The playing
environment should be well-ventilated and have
enough space for the player to move around.
Fabricated
@@
#Knowledge#: 750 Seventh Avenue is a 615 ft (187m)
tall Class-A office skyscraper in New York City. 101
Park Avenue is a 629 ft tall skyscraper in New York City,
New York.
#Question#: 750 7th Avenue and 101 Park Avenue, are
located in which city?
New York City
Faithful
#Knowledge#: 750 Seventh Avenue is a 615 ft (187m)
tall Class-A office skyscraper in New York City. 101
Park Avenue is a 629 ft tall skyscraper in New York City,
New York.
#Question#: 750 7th Avenue and 101 Park Avenue, are
located in which city?
750 7th Avenue and
101 Park Avenue
are located in Albany,
New York
Hallucinated
#Knowledge#: An uncredited Celeste Holm provides the
voice of Addie Ross, the unseen woman who wrote the
titular letter. Celeste Holm (April 29, 1917 - July 15,
2012) was an American stage, film, and television actress.
#Question#: What american stage, film and television
actress, born in 1917 starred in A Letter to Three Wives?
Celeste Holm
Faithful
#Knowledge#: An uncredited Celeste Holm provides the
voice of Addie Ross, the unseen woman who wrote the
titular letter. Celeste Holm (April 29, 1917 - July 15,
2012) was an American stage, film, and television actress.
#Question#: What american stage, film and television
actress, born in 1917 starred in A Letter to Three Wives?
Addie Ross, the
unseen woman who
wrote the titular letter
in A Letter to Three
Wives, was actually
voiced by Meryl Streep.
Hallucinated
@@
What are the names of Donald Duck’s three nephews?
Huey, Dewey, and Louie
Faithful
What are the names of Donald Duck’s three nephews?
Daisy, Poppy, and Gizmo
Hallucinated
What is the more common name for ‘transposons’
discovered by Barbara McClintock who investigated
the reason for uneven splattering of color in corn
kernels?
Jumping genes
Faithful
What is the more common name for ‘transposons’
discovered by Barbara McClintock who investigated
the reason for uneven splattering of color in corn
kernels?
Magic beans
Hallucinated
@@ Appendix H Table for Notations
NN Number of tokens in the generated text $G$
@@
[…]
@@
MKSMKS Model Knowledge Score
ASAS Alignment Score
@@ Appendix I Impact of Validation Set Size on Threshold Values
We investigate how varying the validation set size affects SHINE’s threshold for the NEC dataset and the LLaMA2 model. We compute $\tau^{MKS}$, $\tau_{l}^{AS}$, and $\tau^{AS}_{u}$, while using only pp% of the validation data, varying $p$ from 10% to 100%. The results are shown in Table 11.